I need Help please.

How many milliequivalents of calcium are in a tablespoonful solution containing 250mg of calcium chloride (CaCl 2)? Atomic Weight of Ca++ = 40; Cl- =35.5)
Please Help

Found 2 solutions by venugopalramana, Ms. Tenney:
Answer by venugopalramana(3286)

FIRST FIND MILI GRAM EQUIVALENT MOLECULAR WEIGHT OF CALCIUM CHLORIDE FROM ITS FORMULA ..THERE IS 1 TOM OF CALCIUM AND 2 ATOMS OF CHLORINE IN IT.HENCE M.WT. =1*CA+2*CL=1*40+2*35.5=111...SO
IN 111 MG OF CACL2 WE HAVE 40 MG.EQV. OF CALCIUM
IN 250 MG OF CACL2 WE SHALL HAVE 250*40/111=90 MG.APPROXIMATELY

Answer by Ms. Tenney(1)

first find mili gram equivalent molecular weight.